# Port-Hamiltonian Neural ODE Networks on Lie Groups For Robot Dynamics Learning and Control

### v1

@@ IV-D Reformulation as Port-Hamiltonian Dynamics
To model energy dissipating elements such as friction or drag forces, we reformulate the Hamiltonian dynamics on a matrix Lie group (IV-C) in Port-Hamiltonian form (22). Such elements are often modeled [64] as a linear transformation $\mathbf{D}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})\succeq 0$ of the velocity $\boldsymbol{\xi}$ and only affect the generalized momenta ${\boldsymbol{\mathfrak{p}}}$, i.e.,
@@
The Hamiltonian dynamics (IV-C) is a special case of (22), where the dissipation matrix is $\mathbf{D}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, the input matrix is ${\cal G}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\begin{bmatrix}\bf 0^{\top}&\mathbf{B}({\boldsymbol{\mathfrak{q}}})^{\top}\end{bmatrix}^{\top}$ and the interconnection matrix ${\cal J}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})$ can be obtained by rearranging (IV-C) with $\mathbf{u}=0$ and is guaranteed to be skew-symmetric due to the energy conservation (IV-C).
@@ IV-E Example: Hamiltonian Dynamics on the SE(3) Manifold
The isomorphism between ${\boldsymbol{\mathfrak{s}}\boldsymbol{\mathfrak{e}}}(3)$ and $\mathbb{R}^{6}$ via (11) simplifies the Hamiltonian (IV-C) and its port-Hamiltonian formulation (22) as follows. The Lagrangian function on $SE(3)$ can be expressed in terms of ${\boldsymbol{\mathfrak{q}}}$ and $\boldsymbol{\zeta}$, instead of ${\boldsymbol{\mathfrak{q}}}$ and $\boldsymbol{\xi}$:
@@
and input matrix $\qquad\mathcal{G}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\begin{bmatrix}\bf 0^{\top}&\mathbf{B}({\boldsymbol{\mathfrak{q}}})^{\top}\end{bmatrix}^{\top}$, where ${\boldsymbol{\mathfrak{q}}}^{\times}=\begin{bmatrix}\mathbf{R}^{\top}\!\!\!\!&\bf 0&\bf 0&\bf 0\\ \bf 0&\hat{\mathbf{r}}_{1}^{\top}&\hat{\mathbf{r}}_{2}^{\top}&\hat{\mathbf{r}}_{3}^{\top}\end{bmatrix}^{\top}$. The port-Hamiltonian formulation allows us to model dissipation elements in the dynamics by the dissipation matrix:
@@ V-B Model Architecture
Since robots are physical systems, their dynamics $\mathbf{f}(\mathbf{x},\mathbf{u})$ satisfy the Hamiltonian formulation (Sec. IV-C). To learn the dynamics $\mathbf{f}(\mathbf{x},\mathbf{u})$ from a trajectory dataset ${\cal D}$, we design a neural ODE network (Sec. IV-F), approximating the dynamics via a parametric function $\bar{\mathbf{f}}_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{u})$ based on Eq. (IV-C).
Figure 2: Architecture of port-Hamiltonian neural ODE network on a Lie group. The trainable terms are shown in green.
@@ V-C Training Process
where at each time $t_{k},k=1,\ldots,N$, the adjoint state $\mathbf{a}_{k}$ at time $t_{k}$ is reset to $\frac{\partial L}{\partial\bar{\mathbf{x}}_{k}}$. The resulting derivative $\frac{\partial L}{\partial\boldsymbol{\theta}}$ is used to update the parameters $\boldsymbol{\theta}$ using gradient descent.
@@ V-D Application to SE(3) Hamiltonian Dynamics Learning
When the Hamiltonian dynamics in (IV-C) are defined on the $SE(3)$ manifold, the equations of motion become (IV-E). The neural ODE network architecture in (V-B) is simplified as follows. We use five neural networks with parameters $\boldsymbol{\theta}=(\boldsymbol{\theta}_{\mathbf{v}},\boldsymbol{\theta}_{\boldsymbol{\omega}},\boldsymbol{\theta}_{V},\boldsymbol{\theta}_{\mathbf{D}},\boldsymbol{\theta}_{\mathbf{B}})$ to approximate the blocks $\mathbf{M}^{-1}_{\mathbf{v};{\boldsymbol{\theta}}}({\boldsymbol{\mathfrak{q}}})$, $\mathbf{M}^{-1}_{\boldsymbol{\omega};{\boldsymbol{\theta}}}({\boldsymbol{\mathfrak{q}}})$ of the inverse generalized mass in (25), the potential energy ${\cal V}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}})$, the dissipation matrix $\mathbf{D}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})$, and the input matrix $\mathbf{B}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}})$, respectively. The approximated kinetic energy is calculated as ${\cal T}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\frac{1}{2}{\boldsymbol{\mathfrak{p}}}^{\top}\mathbf{M}^{-1}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}}){\boldsymbol{\mathfrak{p}}}$, where $\mathbf{M}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}})=\operatorname{diag}(\mathbf{M}_{\mathbf{v};{\boldsymbol{\theta}}}({\boldsymbol{\mathfrak{q}}}),\mathbf{M}_{\boldsymbol{\omega};{\boldsymbol{\theta}}}({\boldsymbol{\mathfrak{q}}}))$.
@@ VI Energy-based Control Design
The function $\bar{\mathbf{f}}_{\boldsymbol{\theta}}$ (V-B) learned in Sec. V satisfies the port-Hamiltonian dynamics in Eq. (22) by design. This section extends the interconnection and damping assignment passivity-based control (IDA-PBC) approach [31, 57, 58] to Lie groups to achieve trajectory tracking (Problem 2) based on the learned port-Hamiltonian dynamics. We further derive a tracking controller specifically for learned Hamiltonian dynamics on the $SE(3)$ manifold in Sec. V-D. In the remainder of the paper, we omit the subscript $\boldsymbol{\theta}$ in the learned model for readability.
@@ VI-A IDA-PBC Control Design for Trajectory Tracking
(-(- ( $\mathbf{{\cal J}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})-\mathbf{{\cal R}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}}))\begin{bmatrix}\frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{q}}}}\\ \frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}+\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}\\ \dot{{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}-\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}_{e}\\ \dot{{\boldsymbol{\mathfrak{p}}}}_{e}\end{bmatrix}.$ (57)
@@
$\mathbf{u}_{ES}$ $=\mathbf{{\cal G}}^{\dagger}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})\left((\mathcal{J}_{d}({\boldsymbol{\mathfrak{q}}}_{e},{\boldsymbol{\mathfrak{p}}}_{e}))\begin{bmatrix}\frac{\partial\mathcal{H}_{d}}{\partial{\boldsymbol{\mathfrak{q}}}_{e}}\\ \frac{\partial\mathcal{H}_{d}}{\partial{\boldsymbol{\mathfrak{p}}}_{e}}\end{bmatrix}-\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}_{e}\\ \dot{{\boldsymbol{\mathfrak{p}}}}_{e}\end{bmatrix}\right.$ (58e)
$\left.\qquad-(\mathbf{{\cal J}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})-\mathbf{{\cal R}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}}))\begin{bmatrix}\frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{q}}}}\\ \frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}+\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}\\ \dot{{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}\right),$ (58j)
$\mathbf{u}_{DI}$ $=-\mathbf{{\cal G}}^{\dagger}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})\mathcal{R}_{d}({\boldsymbol{\mathfrak{q}}}_{e},{\boldsymbol{\mathfrak{p}}}_{e})\begin{bmatrix}\frac{\partial\mathcal{H}_{d}}{\partial{\boldsymbol{\mathfrak{q}}}_{e}}\\ \frac{\partial\mathcal{H}_{d}}{\partial{\boldsymbol{\mathfrak{p}}}_{e}}\end{bmatrix},$ (58m)
@@
-- $\left.({\cal J}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})-{\cal R}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}}))\begin{bmatrix}\frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{q}}}}\\ \frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}+\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}\\ \dot{{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}-\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}_{e}\\ \dot{{\boldsymbol{\mathfrak{p}}}}_{e}\end{bmatrix}\right)=0.$ (68)
@@ VI-B Tracking Control for Port-Hamiltonian Dynamics on the SE(3) Manifold
where $\mathbf{M}_{d}({\boldsymbol{\mathfrak{q}}}_{e})$ and $V_{d}({\boldsymbol{\mathfrak{q}}}_{e})$ are the desired generalized mass and potential energy. Choosing the following desired inter-connection matrix and dissipation matrix:
@@
since $\dot{\mathbf{R}}_{e}=\frac{d}{dt}(\mathbf{R}_{e})=\mathbf{R}_{e}\widehat{\boldsymbol{\omega}}_{e}$ as shown in [70, Sec. III-A] and $\dot{\mathbf{p}}-\dot{\mathbf{p}}^{*}~{}=~{}\mathbf{R}\mathbf{v}-\mathbf{R}^{*}\mathbf{v}^{*}=\mathbf{R}\mathbf{v}_{e}$.
@@
Consider a port-Hamiltonian system on the $SE(3)$ manifold with dynamics (IV-E). Assume that the matching condition (76) is satisfied, the desired momentum’s derivative $\dot{{\boldsymbol{\mathfrak{p}}}}^{*}$ is bounded, and the matrices $\mathbf{K}_{\mathbf{p}}$, $\mathbf{K}_{\mathbf{R}}$, and $\mathbf{K}_{\mathbf{d}}$ are positive-definite. The control policy in (VI-B) leads to closed-loop error dynamics in (47), (71). The tracking error $({\boldsymbol{\mathfrak{q}}}_{e},{\boldsymbol{\mathfrak{p}}}_{e})=\left((\mathbf{p}_{e},\mathbf{R}_{e}),{\boldsymbol{\mathfrak{p}}}_{e}\right)$ asymptotically stabilizes to $\left((\bf 0,\mathbf{I}),\bf 0\right)$ with Lyapunov function given by the desired Hamiltonian ${\cal H}_{d}({\boldsymbol{\mathfrak{q}}}_{e},{\boldsymbol{\mathfrak{p}}}_{e})$ in (77).
@@
Since the matching condition is satisfied and the desired momentum’s derivative $\dot{{\boldsymbol{\mathfrak{p}}}}^{*}$ is bounded, the control policy in (VI-B), (75b) exists and achieves the desired closed-loop error dynamics:
@@
Without requiring a priori knowledge of the system parameters, the control design in (VI-B) offers a unified control approach for $SE(3)$ Hamiltonian systems that achieves trajectory tracking, if permissible by the system’s degree of underactuation. Thus, our control design solves Problem 2 for rigid-body robot systems, such as UGVs, UAVs, and UUVs, with tracking performance guaranteed by Theorem 1.

### v2

@@ IV-D Reformulation as Port-Hamiltonian Dynamics
To model energy dissipating elements such as friction or drag forces, we reformulate the Hamiltonian dynamics on a matrix Lie group (18) in port-Hamiltonian form (22). Such elements are often modeled [70] as a linear transformation $\mathbf{D}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})\succeq 0$ of the velocity $\boldsymbol{\xi}$ and only affect the generalized momenta ${\boldsymbol{\mathfrak{p}}}$, i.e.,
@@
The Hamiltonian dynamics (18) is a special case of (22), where the dissipation matrix is $\mathbf{D}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\bf 0$, the input matrix is ${\cal G}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\begin{bmatrix}\bf 0^{\top}&\mathbf{B}({\boldsymbol{\mathfrak{q}}})^{\top}\end{bmatrix}^{\top}$ and the interconnection matrix ${\cal J}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})$ can be obtained by rearranging (18) with $\mathbf{u}=0$ and is guaranteed to be skew-symmetric due to the energy conservation (IV-C). In an implementation, the coordinates ${\boldsymbol{\mathfrak{q}}}$ and momentum ${\boldsymbol{\mathfrak{p}}}$ may be represented as vectors in $\mathbb{R}^{n^{2}}$, leading to an interconnection matrix ${\cal J}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})\in\mathbb{R}^{2n^{2}\times 2n^{2}}$.
@@ IV-E Example: Hamiltonian Dynamics on the SE(3) Manifold
The isomorphism between ${\boldsymbol{\mathfrak{s}}\boldsymbol{\mathfrak{e}}}(3)$ and $\mathbb{R}^{6}$ via (12) simplifies the Hamiltonian (18) and its port-Hamiltonian formulation (22) as follows. The Lagrangian function on $SE(3)$ can be expressed in terms of ${\boldsymbol{\mathfrak{q}}}$ and $\boldsymbol{\zeta}$, instead of ${\boldsymbol{\mathfrak{q}}}$ and $\boldsymbol{\xi}$:
@@
and input matrix $\qquad\mathcal{G}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\begin{bmatrix}\bf 0^{\top}&\mathbf{B}({\boldsymbol{\mathfrak{q}}})^{\top}\end{bmatrix}^{\top}$, where ${\boldsymbol{\mathfrak{q}}}^{\times}=\begin{bmatrix}\mathbf{R}^{\top}\!\!\!\!&\bf 0&\bf 0&\bf 0\\ \bf 0&\hat{\mathbf{r}}_{1}^{\top}&\hat{\mathbf{r}}_{2}^{\top}&\hat{\mathbf{r}}_{3}^{\top}\end{bmatrix}^{\top}$. Note that the kinematics constraints $\dot{{\boldsymbol{\mathfrak{q}}}}={\boldsymbol{\mathfrak{q}}}\boldsymbol{\xi}$ of the coordinates ${\boldsymbol{\mathfrak{q}}}$ on $SE(3)$, which guarantee that the coordinates ${\boldsymbol{\mathfrak{q}}}$ remain in $SE(3)$, are not affected by vectorization. The port-Hamiltonian formulation allows us to model dissipation elements in the dynamics by the dissipation matrix:
@@ IV-F Neural ODE Networks
Recently, neural ODE networks have been extended from Euclidean space to Lie groups [19, 20, 21, 22], guaranteeing that the Lie group constraints are satisfied by the predicted states by design. While it is possible to train our Hamiltonian dynamics model using a Lie group neural ODE network, we leave this investigation for future work due to the lack of suitable open-source software for Lie group ODE integration.
@@ V-B Model Architecture
Since robots are physical systems, their dynamics $\mathbf{f}(\mathbf{x},\mathbf{u})$ satisfy the Hamiltonian formulation (Sec. IV-C). To learn the dynamics $\mathbf{f}(\mathbf{x},\mathbf{u})$ from a trajectory dataset ${\cal D}$, we design a neural ODE network (Sec. IV-F), approximating the dynamics via a parametric function $\bar{\mathbf{f}}_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{u})$ based on Eq. (18).
Figure 2: Architecture of port-Hamiltonian neural ODE network on matrix Lie group. The trainable terms are shown in green.
@@ V-C Training Process
where at each time $t_{k},k=1,\ldots,N$, the adjoint state $\mathbf{a}_{k}$ at time $t_{k}$ is reset to $\frac{\partial L}{\partial\bar{\mathbf{x}}_{k}}$. The resulting derivative $\frac{\partial L}{\partial\boldsymbol{\theta}}$ is used to update the parameters $\boldsymbol{\theta}$ using gradient descent. Note that even though the Lie group and Hamiltonian structures are preserved in the continuous-time dynamics function $\bar{\mathbf{f}}$, the Lie group constraints might still be violated by the predicted states when the model is trained with a neural ODE network on the embedding space $\mathbb{R}^{n\times n}$ instead of on the matrix Lie group [21, 22]. To prevent this, we use a high-order integrator, such as $5$th-order Runge-Kutta [76], in the Euclidean space neural ODE, for which open-source software is available [31, 77]. Investigating how to train Hamiltonian models with a neural ODE network defined directly on the Lie group [78, 21, 22] is an interesting future direction. For example, the Lie group neural ODE by Wotte et al. [22] offers an approach to guarantee Lie group constraints in the dynamics model. Closely related to this direction, Duruisseaux et al. [54] learn discrete-time Hamiltonian dynamics while guaranteeing Lie group constraints by design using variational integration.
@@ V-D Application to SE(3) Hamiltonian Dynamics Learning
When the Hamiltonian dynamics in (18) are defined on the $SE(3)$ manifold, the equations of motion become (30). The neural ODE network architecture in (33) is simplified as follows. We use five neural networks with parameters $\boldsymbol{\theta}=(\boldsymbol{\theta}_{\mathbf{v}},\boldsymbol{\theta}_{\boldsymbol{\omega}},\boldsymbol{\theta}_{V},\boldsymbol{\theta}_{\mathbf{D}},\boldsymbol{\theta}_{\mathbf{B}})$ to approximate the blocks $\mathbf{M}^{-1}_{\mathbf{v};{\boldsymbol{\theta}}}({\boldsymbol{\mathfrak{q}}})$, $\mathbf{M}^{-1}_{\boldsymbol{\omega};{\boldsymbol{\theta}}}({\boldsymbol{\mathfrak{q}}})$ of the inverse generalized mass in (25), the potential energy ${\cal V}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}})$, the dissipation matrix $\mathbf{D}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})$, and the input matrix $\mathbf{B}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}})$, respectively. The approximated kinetic energy is calculated as ${\cal T}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})=\frac{1}{2}{\boldsymbol{\mathfrak{p}}}^{\top}\mathbf{M}^{-1}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}}){\boldsymbol{\mathfrak{p}}}$, where $\mathbf{M}_{\boldsymbol{\theta}}({\boldsymbol{\mathfrak{q}}})=\operatorname{diag}(\mathbf{M}_{\mathbf{v};{\boldsymbol{\theta}}}({\boldsymbol{\mathfrak{q}}}),\mathbf{M}_{\boldsymbol{\omega};{\boldsymbol{\theta}}}({\boldsymbol{\mathfrak{q}}}))$.
@@ VI Energy-based Control Design
The function $\bar{\mathbf{f}}_{\boldsymbol{\theta}}$ (33) learned in Sec. V satisfies the port-Hamiltonian dynamics in Eq. (22) by design. This section extends the interconnection and damping assignment passivity-based control (IDA-PBC) approach [51, 61, 62] to Lie groups to achieve trajectory tracking (Problem 2) based on the learned port-Hamiltonian dynamics. We further derive a tracking controller specifically for learned Hamiltonian dynamics on the $SE(3)$ manifold in Sec. V-D. In the remainder of the paper, we omit the subscript $\boldsymbol{\theta}$ in the learned model for readability.
@@ VI-A IDA-PBC Control Design for Trajectory Tracking
(-(- ( $\mathbf{{\cal J}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})-\mathbf{{\cal R}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}}))\begin{bmatrix}\frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{q}}}}\\ \frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}+\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}\\ \dot{{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}-\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}_{e}\\ \dot{{\boldsymbol{\mathfrak{p}}}}_{e}\end{bmatrix}.$
@@
$\mathbf{u}_{ES}$ $=\mathbf{{\cal G}}^{\dagger}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})\left((\mathcal{J}_{d}({\boldsymbol{\mathfrak{q}}}_{e},{\boldsymbol{\mathfrak{p}}}_{e}))\begin{bmatrix}\frac{\partial\mathcal{H}_{d}}{\partial{\boldsymbol{\mathfrak{q}}}_{e}}\\ \frac{\partial\mathcal{H}_{d}}{\partial{\boldsymbol{\mathfrak{p}}}_{e}}\end{bmatrix}-\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}_{e}\\ \dot{{\boldsymbol{\mathfrak{p}}}}_{e}\end{bmatrix}\right.$ (49a)
$\left.\qquad-(\mathbf{{\cal J}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})-\mathbf{{\cal R}}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}}))\begin{bmatrix}\frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{q}}}}\\ \frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}+\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}\\ \dot{{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}\right),$
$\mathbf{u}_{DI}$ $=-\mathbf{{\cal G}}^{\dagger}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})\mathcal{R}_{d}({\boldsymbol{\mathfrak{q}}}_{e},{\boldsymbol{\mathfrak{p}}}_{e})\begin{bmatrix}\frac{\partial\mathcal{H}_{d}}{\partial{\boldsymbol{\mathfrak{q}}}_{e}}\\ \frac{\partial\mathcal{H}_{d}}{\partial{\boldsymbol{\mathfrak{p}}}_{e}}\end{bmatrix},$ (49b)
@@
-- $\left.({\cal J}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}})-{\cal R}({\boldsymbol{\mathfrak{q}}},{\boldsymbol{\mathfrak{p}}}))\begin{bmatrix}\frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{q}}}}\\ \frac{\partial\mathcal{H}}{\partial{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}+\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}\\ \dot{{\boldsymbol{\mathfrak{p}}}}\end{bmatrix}-\begin{bmatrix}\dot{{\boldsymbol{\mathfrak{q}}}}_{e}\\ \dot{{\boldsymbol{\mathfrak{p}}}}_{e}\end{bmatrix}\right)=0.$
@@ VI-B Tracking Control for Port-Hamiltonian Dynamics on the SE(3) Manifold
where $\mathbf{M}_{d}({\boldsymbol{\mathfrak{q}}}_{e})$ and $V_{d}({\boldsymbol{\mathfrak{q}}}_{e})$ are the desired generalized mass and potential energy.
Choosing the following desired inter-connection matrix and dissipation matrix:
@@
since $\dot{\mathbf{R}}_{e}=\frac{d}{dt}(\mathbf{R}_{e})=\mathbf{R}_{e}\widehat{\boldsymbol{\omega}}_{e}$ as shown in [79, Sec. III-A] and $\dot{\mathbf{p}}-\dot{\mathbf{p}}^{*}~{}=~{}\mathbf{R}\mathbf{v}-\mathbf{R}^{*}\mathbf{v}^{*}=\mathbf{R}\mathbf{v}_{e}$.
@@
Consider a port-Hamiltonian system on the $SE(3)$ manifold with dynamics (30). Assume that the matching condition (58) is satisfied, the desired momentum’s derivative $\dot{{\boldsymbol{\mathfrak{p}}}}^{*}$ is bounded, and the matrices $\mathbf{K}_{\mathbf{p}}$, $\mathbf{K}_{\mathbf{R}}$, and $\mathbf{K}_{\mathbf{d}}$ are positive-definite. The control policy in (57) leads to closed-loop error dynamics in (47), (53). The tracking error $({\boldsymbol{\mathfrak{q}}}_{e},{\boldsymbol{\mathfrak{p}}}_{e})=\left((\mathbf{p}_{e},\mathbf{R}_{e}),{\boldsymbol{\mathfrak{p}}}_{e}\right)$ asymptotically stabilizes to $\left((\bf 0,\mathbf{I}),\bf 0\right)$ with Lyapunov function given by the desired Hamiltonian ${\cal H}_{d}({\boldsymbol{\mathfrak{q}}}_{e},{\boldsymbol{\mathfrak{p}}}_{e})$ in (59).
@@
Since the matching condition is satisfied and the desired momentum’s derivative $\dot{{\boldsymbol{\mathfrak{p}}}}^{*}$ is bounded, the control policy in (57), (57b) exists and achieves the desired closed-loop error dynamics:
@@
Without requiring a priori knowledge of the system parameters, the control design in (57) offers a unified control approach for $SE(3)$ Hamiltonian systems that achieves trajectory tracking, if permissible by the system’s degree of underactuation. Thus, our control design solves Problem 2 for rigid-body robot systems, such as UGVs, UAVs, and UUVs, with tracking performance guaranteed by Theorem 1.